OpenStudy (anonymous):

(1/3)^50 : (-9)^25 - 2/3 : 4

OpenStudy (anonymous):

@Jhannybean @Hero .... CAN U HELP ME ? I AM HURRY

OpenStudy (anonymous):

@Michele_Laino @Mimi_x3

OpenStudy (anonymous):

@Ganpat can u help me ?

OpenStudy (ganpat):

\[(1/3)^{50} : (-9)^{25} = (3)^{-50} : [(-3)^{25} * (3)^{25}] = (3)^{-75} : (-3)^{25}\] I highly doubt my solution, but its worth a try, Allow others to confirm over it.

OpenStudy (zzr0ck3r):

Why are you in a hurry?

OpenStudy (ganpat):

@zzr0ck3r : me ?? No man, i have whole day with me. :D btw could you please help ??

OpenStudy (zzr0ck3r):

Not you

OpenStudy (anonymous):

me ?

OpenStudy (zzr0ck3r):

the person that said they were in a hurry lol

OpenStudy (zzr0ck3r):

I don't know what the question is

OpenStudy (anonymous):

the question is compute it...

OpenStudy (zzr0ck3r):

ahh this is out of my world. this looks like Chinese to me.

OpenStudy (zzr0ck3r):

this is definitatly not ture \((-9)^{25}=3^{-50}\)

OpenStudy (zzr0ck3r):

one is positive one is negative

OpenStudy (ganpat):

i agree, So i tried breaking it in two bases 3^25 and -3^25

OpenStudy (zzr0ck3r):

I am not sure what you are doing but \((-9)^{25}=-(9)^{25}=-(3^2)^{25}=-(3^{50})=-(3^{25})(3^{25})\)

OpenStudy (anonymous):

yes . right @zzr0ck3r

OpenStudy (ganpat):

(−9)^25=−(9)^25 This mean whatever will be the solution for 9^25 has to be negative. I doubt this thing.

OpenStudy (ganpat):

\[(-9)^{25} = -(9)^{25}\] This mean whatever will be the solution for 9^25 has to be negative, I doubt this thing.

OpenStudy (anonymous):